OpenStudy (anonymous):

Pre algebra , You have a 5-question multiple-choice test.each question has 4 choices .you dont know any of the answers.what is the experimental probability that you will guess exactly 3 out of five questions correctly ? type your sentences below using complete sentences . HELP !!

OpenStudy (anonymous):

you have a \(\frac{1}{4}\) chance of getting any one question right and so a \(\frac{3}{4}\) chance of getting it wrong to get exactly 3 out of 5 right, you get 3 right and 2 wrong there are \(\binom{5}{2}=\frac{5\times 4}{2}=10\) different ways to get 3 out of 5 right, and the probability of each is \((\frac{1}{4})^3\times (\frac{3}{4})^2\) so the probability of getting exactly 3 out of 5 right is \[10\times (\frac{1}{4})^3\times (\frac{3}{4})^2\]

OpenStudy (anonymous):

omg thank you so much i been stuck on this forever

OpenStudy (anonymous):

yw i would not call this pre algebra, or even algebra it is probability these are called bernoulli trials, and the idea is not for a pre algebra class

OpenStudy (anonymous):

oops... :$ heh